• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name D. Lhuillier

    Log entry time 21:51:43 on July20,2005

    Entry number 148777

    keyword=Compton tune

    New beam tune reduced the photon background in the compton detector. S/B is now 4 instead of 2. Only changes in the hall A beam line were the settings of the two quads upstream and dowstream the compton chicane: +15% on 1C20 and -20% on 1H01.
    Background in the electron detector is worse (S/B=1) and still shows some beam halo or energy tail. Nevertheless accurate analysis relies more on photon detector.
    Now fine tuning the apparatus for production mode and systematic studies.
    So far polarization measurements are stable at 86% within stat error (few hours of data).